OpenStudy (anonymous):

Find the value of c: -8.4 = c - 3.1.

OpenStudy (anonymous):

c = -5.3 c = -11.5 c = 5.3 c = 11.5 answers

OpenStudy (rsadhvika):

add 3.1 both sides

OpenStudy (nurali):

-8.4 = c - 3.1 c=-8.4+3.1 c=-5.3